:>roblem 3-27A R ecordingevents in a statements model and I-accounts a preparing a trial balance f he following acco unting eve nts a pply to Parks Co. for the yea r 201 1: !\sset Source Transactions

I. Began o pera ti o ns when th e business acq uired S 20.000 cash from th e iss ue stock. 2. Performed services and co llecte d cash of S, 1 000. 3. Co llec te d $4,500 of cash in advance for services to be provided over th e next 12 4. Provided SI 2.000 of service s o n account. 5. Purchased supp lies of $420 on acco unt. Asset Exchange Transactions 6. Pur c hased S4.000 of land for cash. 7. Collec ted S8.500 of cas h from account s receivable. 8. Purchased $500 of supp lies wi th cas h. 9. Pa id S3.600 in advance for one year's rent.
Answer #1
3-27A)
Recording events in a statements model
ie. Accounting Equation
Event Assets = Liabilities + Stockholders'Equity
Cash Acc Rec. Supplies Prepaid Rent Land = Unearned Rev. Acc payable Comm Stock S.Revenue
1 20000 = 20000
2 1000 = 1000
3 4500 = 4500
4 12000 = 12000
5 420 = 420
6 -4000 4000 =
7 8500 -8500 =
8 -500 500 =
9 -3600                                 3600                    =                                                                                            
Balance 25900 3500 920 3600 4000 = 4500 420 20000 13000
G Total 37920 = 37920
Trial Balance:
Acc. Title Debit $ Credit $
Cash 25900
Acc Rec. 3500
Supplies 920
Prepaid rent 3600
land 4000
Unearned revenue 4500
Account Payable 420
Common Stock 20000
Service Revenue               13000
Total 37920 37920